On April 5, 2019, Ryan received land and a building from Thom as a gift. Thom's adjusted basis and the fair market value at the date of the gift are as follows:

Asset

Adjusted Basis

FMV

Land

$40,000

53,000

Building

110,000

120,000

Thom paid gift tax of $8,000 on the transfer.

When required, round any division to four decimal places. Round final answers to the nearest dollar.

Determine Ryan's adjusted basis for the land and building:

1) Ryan's total basis for the assets received from Thom is _______.

2) Ryan's adjusted basis for the land is _______

3) Ryan's adjusted basis for the building is ________
Answer #1

Part 1

Ryan's total basis for the assets received from Thom = $40000+$110000 = $150000

Part 2

Ryan's adjusted basis for the land = (53000/173000)*150000 = $45954

Part 3

Ryan's adjusted basis for the building = (120000/173000)*150000 = $104046